LABORATORY DETECTION OF TWELVE CARBON-SULFUR CHAINS

\underline{V. D. GORDON}, M. C. McCARTHY, A. J. APPONI, and P. THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138 and Division of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138.

Twelve sulfur-containing carbon chains have recently been detected and spectroscopically characterized in our laboratory by Fourier transform microwave spectroscopy. These include: the singlet chains C_{7}S and C_{9}S; the triplet chains C_{6}S and C_{8}S; the free radicals HC_{5}S, HC_{6}S, HC_{7}S, and HC_{8}S; and the asymmetric tops H_{2}C_{4}S, H_{2}C_{5}S, H_{2}C_{6}S, and H_{2}C_{7}S. In addition, all three fine structure ladders of triplet C_{4}S have now been observed, and the two fine-structure constants γ and λ determined to high accuracy. An experimental structure of C_{5}S has also been derived on the basis of the singlysubstituted isotopic species which were observed in natural abundance. A summary of these and other recent results will be presented.